Fill in the blanks with >,
 (–21) – (–10) _____ (–31) + (–11)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

-21–(-10) = -21 + 10 = -11
-31 + (-11) = -42
-11 > -42
Therefore (-21)–(-10) > (-31) + (-11)

Thanks